find the pair of consecutive integers such that the sum of twice the smaller even integer and 10 times the larger even integer is 500
Found 2 solutions by ReadingBoosters, mananth:
Answer by ReadingBoosters(3246)

Given
x x+2
2x+10(x+2)=500
...
Solve for x
..
2x + 10x + 20 = 500
12x = 500 - 20
12x = 480
x = 40
...
Integers are 40 and 42
...
Check

....................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com

Answer by mananth(16946)

let the even integers be x & x+2
2x+10(x+2)=500
2x+10x+20=500
30x=500-20
30x=480
/30
x= 16
The consecutive even integers are 16 & 18
m.ananth@hotmail.ca